# DCT-MARL: A Dynamic Communication Topology Based MARL Algorithm for Platoon Control

### v1

@@ Abstract
With the rapid advancement of vehicular communication and autonomous driving technologies, connected vehicle platoon has emerged as a promising approach to improve traffic efficiency and driving safety. Reliable Vehicle-to-Vehicle (V2V) communication is critical to achieving efficient cooperative control. However, in real-world traffic environments, V2V links may suffer from time-varying delay and packet loss, leading to degraded control performance and even safety risks. To mitigate the adverse effects of non-ideal communication, this paper proposes a Dynamic Communication Topology based Multi-Agent Reinforcement Learning (DCT-MARL) algorithm for robust cooperative platoon control. Specifically, the state space is augmented with historical control action and delay to enhance robustness against communication delay. To mitigate the impact of packet loss, a multi-key gated communication mechanism is introduced, which dynamically adjusts the communication topology based on the correlation between agents and their current communication status. Simulation results demonstrate that the proposed DCT-MARL significantly outperforms state-of-the-art methods in terms of string stability and driving comfort, validating its superior robustness and effectiveness.
@@ I Introduction
In recent years, with the rapid deployment of Vehicle-to-Vehicle (V2V) communication facilities and autonomous driving technologies, connected vehicle platoon has become a promising approach for enhancing road safety and traffic efficiency [1]. Through V2V Communication, connected vehicles can form self-organized platoons and maintain a short inter-vehicle distance and coordinated driving behavior, thereby improving road throughput while reducing energy consumption and emissions [2].
@@
In such platoon systems, the effectiveness of cooperative control critically depends on the real-time sharing of vehicle driving state information (e.g., position, velocity, and acceleration, etc.). Consequently, reliable V2V communication is a essential prerequisite for ensuring control performance [3]. However, in real-world traffic environments, channel fading, signal obstruction, and spectral interference may induce substantial communication quality variations, resulting in time-varying delay and packet loss.
@@
These imperfections not only hinder timely and accurate information exchange, as delay leads to outdated data and packet loss results in missing information, but also directly affect the velocity stability and driving comfort of the platoon system [4]. Hence, developing robust platoon control strategies under non-ideal communication conditions has become a pressing research challenge.
@@
To mitigate delay effects, a range of delay compensation based robust control strategies have been proposed. Some studies adopt deterministic or random delay modeling, allowing controllers to compensate for delay-induced performance degradation [5, 6]. More recently, researchers have also explored solutions within the Reinforcement Learning (RL) framework by expanding the agent’s state space to include historical control actions[7, 8], which improves the system’s robustness under delayed communication. Nevertheless, these methods largely rely on simplified delay assumptions and lack explicit sensitivity to the time-varying nature of delay, limiting their scalability to highly dynamic communication environments.
@@
For packet loss, mitigation strategies typically rely on holding previous inputs, interpolating historical data, or fallback modes [9, 10, 11, 12, 13, 14]. Although these methods can maintain basic control performance under mild packet loss, they often fail to guarantee stability and safety in high-loss or highly dynamic environments. Some recent studies have introduced dynamic communication topology mechanisms that adjust weights based on communication quality, aiming to mitigate performance degradation caused by communication disruptions [15, 16]. While this mitigates performance degradation from communication packet loss, these strategies focus solely on communication quality metrics and overlook the need to consider the relevance of communication vehicles in the current dynamic traffic environment when adjusting the communication topology. For example, when a human driver accelerates, they pay attention to the vehicle ahead, while when decelerating, they focus on the vehicle behind[17].
@@
Moreover, most existing studies typically address either communication delay or packet loss in isolation, lacking a unified control framework that considers their coupled impact. Although some traditional optimization-based works have proposed integrated modeling approaches [9, 18, 5, 19, 20], they typically rely on accurate system models and analytical formulations, which limits their adaptability and generalization under real-world uncertainties and unmodeled dynamics. Therefore, it’s necessary to investigate novel intelligent algorithms to solve the adaptability problem in dynamic and complex traffic scenarios.
@@
In contrast, Multi-Agent Reinforcement Learning (MARL), has recently emerged as a promising solution for cooperative platoon control due to its model-free nature, strong generalization capabilities, and support for distributed decision-making [21, 22, 23]. Under assumptions of ideal communication or deterministic delay, previous studies have validated MARL’s efficacy in platoon control [24, 25]. Some methods further enhance MARL by introducing communication-aware designs, that though consider inter-agent information correlations to build dynamic communication topologies [26]. However, these approaches tend to focus on correlation among agents, while overlooking the dynamic of communication quality.
@@
To bridge this gap, this paper proposes a novel Dynamic Communication Topology based Multi-Agent Reinforcement Learning (DCT-MARL) algorithm to enhance the control performance of connected vehicle platoon under non-ideal communication conditions. By dynamically adjusting the communication topology and augmenting the state space, the proposed method improves the robustness and adaptability of vehicle platoon in realistic communication environments. The main contributions are summarized as follows:
@@
    A multi-key gated communication network is proposed, enabling each agent to infer the relevance of communication agents based on real-time observations and historical context. This mechanism dynamically adjusts the communication topology in response to packet loss.
@@
    A state augmentation scheme is designed to mitigate the impact of communication delay on platoon control. Specifically, the agent’s state is extended to include both the history control action and delay information. This augmentation enhances the policy’s sensitivity to time-varying communication dealys and improves the system’s robustness and adaptability under delayed communication.
@@
    A co-simulation environment integrating vehicular control and wireless communication is developed to enable comprehensive evaluation under realistic conditions. Extensive experiments demonstrate that the proposed DCT-MARL algorithm outperforms state-of-the-art methods in terms of string stability and driving comfort under non-ideal communication conditions.
@@ II-A Platoon Control under Unideal Communication Conditions
Reliable V2V communication is essential for cooperative platoon control, however, packet loss and delays in practical V2V communication may induce performance degradation in the control system[4, 3].
@@
To mitigate the negative impact of communication delay, a range of delay compensation based robust control strategies have been proposed[5, 6, 7, 8, 27]. To address packet loss, traditional approaches employ techniques include Zero-Order Hold (ZOH) to maintain previous control inputs [9], historical data interpolation [10, 11], or fallback modes such as Adaptive Cruise Control (ACC) [12, 13, 14]. Although these methods can maintain basic control feasibility under mild packet loss, they often fail to ensure stability in scenarios with the phenomenon of fluctuating velocities and persistent packet loss. Moreover, predictive strategies may impose substantial computational burdens. To enhance adaptability, some recent studies have introduced dynamic communication topology mechanisms that adjust weights based on communication quality, thereby mitigating the effects of link failures [15, 16]. For example, Liu et al. [15] developed an traditional optimization-based method to adjust weights according to communication conditions, whereas Wang et al. [16] proposed a self-adaptive topology control method based on a Dual-Decision Deterministic Policy Gradient ($D^{3}$PG) algorithm for platoon control, which significantly enhances system resilience under fluctuating communication conditions. However, these strategies primarily focus on communication quality metrics, overlooking the fact that in practical applications, the selection of communication targets should be context-aware and responsive to dynamic traffic conditions.
@@
Furthermore, most existing works consider either delay or packet loss in isolation, and are thus inadequate for scenarios where both coexist. In practice, delay and packet loss often appear simultaneously, imposing compounded challenges. To address this, some recent works have proposed unified models that jointly consider the effects of delay and packet loss[9, 18, 5, 19, 20]. For instance, Vegamoor et al.[18] modeled delay and packet loss simultaneously, and introduced a safety-aware constraint on distance errors between any two vehicles and derived a lower bound on safe time gaps to ensure string stability under unideal communication scenarios. Zhao et al.[5] investigated control strategies under constant time headway distance, modeling non-ideal conditions including limited communication range, random packet loss, and time-varying delay. Their approach employed a distance-based dynamic topology and a robust controller that utilized multi-leader an follower information. Following this, Wang et al.[20] further refined the approach by proposing a CAV following model based on dynamic weight optimization, which adaptively adjusts the importance of incoming information. Their framework incorporated realistic vehicle dynamics and demonstrated strong robustness under complex and adverse communication conditions.
@@
While these control approaches provide valuable theoretical foundations, they often depend on accurate system modeling and incur significant computational overhead. As a result, they are difficult to scale in dynamic and uncertain traffic environments characterized by rapidly changing velocities and nonlinear interactions.
@@ II-B MARL-based Platoon Control
In recent years, MARL has gained significant attention in intelligent transportation systems due to its strengths in distributed decision-making, policy learning, and environmental adaptability[24, 28, 29, 25]. MARL has become a promising paradigm for cooperative control tasks, such as vehicle platoon. For example, Chen et al.[28] evaluated MARL frameworks using both local and global reward structures in platoons of three and five vehicles, and found that decentralized learning with local rewards achieved superior performance in terms of stability and responsiveness compared to centralized approaches.
However, the vast majority of existing MARL-based platoon control studies assume ideal V2V communication, free of delay and packet loss, and thus lack a comprehensive evaluation into the robustness of learned policies under realistic communication impairments. To address this gap, Wang et al. [25] proposed a delay-aware MARL framework with a fixed communication topology, formulated within the partially observable Markov game (POMG), capturing delay-induced uncertainties in platoon systems. Their simulations demonstrated improved platoon stability under time-varying delay. However, existing works still fall short of simultaneously addressing communication delay and packet loss in a unified MARL framework.
On dynamic communication topologies within MARL, researchers have explored inter-agent information correlations to construct adaptive communication topologies. Seminal algorithms including ITGCNet[30], Gated-ACML[31], I2C[26], and NeurComm[32]. In particular, Ding et al. [26] propose I2C, a dynamic communication framework where each agent learns to infer its communication targets from local observations via a prior network trained through causal inference. However, these methods primarily emphasize inter-agent information relevance while neglecting the dynamic variability of the communication environment, limiting their applicability under real-world communication environment.
To overcome these limitations, this paper proposes a DCT-MARL framework that jointly addresses communication delay and packet loss. The framework aims to substantially enhance the robustness and adaptability of connected vehicle platoons in non-ideal communication environment.
@@ III-A Vehicle Dynamics Model
This paper considers a platoon consisting of $N+1$ vehicles, indexed as $0,1,\ldots,n$, where the leading vehicle is indexed as 0, and the following vehicles are indexed from $1$ to $n$. All vehicles communicate through V2V wireless communication to exchange information and coordinate movements. This paper primarily focus on a linear model of a connected automated vehicle (CAV) system, which has been extensively used in vehicle control. The vehicle dynamics are given by [33].
@@
This paper adopt the constant time headway policy (CTHP), meaning that the platoon maintains a constant time gap between vehicles. Each following vehicle $i$ aims to maintain a desired headway distance from its preceding vehicle i1i-1, which is defined as the bumper-to-bumper distance. The ideal inter-vehicle distance between vehicle $i$ and its predecessor i1i-1 is denoted as $d_{r,i,t}$.
@@
Since all vehicles in the platoon are assumed to be identical, this paper consider $d_{0}$ and $l$ to be uniform across all vehicles and omit them in subsequent derivations.
For any vehicle pair $p_{(}i,j)$ in the communication graph, the tracking error consists of both distance error and velocity error, defined as follows:
@@
[…]
where $j<i$ (i.e., vehicle $j$ is ahead of vehicle $i$), the distance error between vehicles $i$ and $j$ is the difference between the actual position minus the velocity of all vehicles from $j+1$ to $i$ multiplied by the sum of $h$.
Each vehicle $i$ is capable of communicating with other vehicle $j$ within its communication range. However, due to potential communication disruptions or packet loss, the communication link may become unavailable at certain time steps. Specifically, if $(i,j)\notin E_{t}$ at certain time steps, vehicle $i$ fails to obtain information from vehicle $j$.
Additionally, due to network congestion, signal attenuation, or interference, communication may experience a delay $\xi_{i,j}$, meaning that vehicle $i$ at time $t$ receives information from vehicle $j$ that corresponds to time $t-\displaystyle\xi_{i,j}$ instead of real-time data.
@@ III-B Control Objectives of Vehicle Platoon
where $\mathrm{acc}_{i,t}$ is the actual acceleration of vehicle $i$, and ui,tu{i,t} is the control input (i.e., the commanded acceleration). The bounds $\mathrm{acc}_{\min}$ and $\mathrm{acc}_{\max}$ represent the allowable acceleration range. These constraints ensure safe and comfortable driving and prevent excessive control efforts.

### v2

@@ Abstract
With the rapid advancement of vehicular communication facilities and autonomous driving technologies, connected vehicle platooning has emerged as a promising approach to improve traffic efficiency and driving safety. Reliable Vehicle-to-Vehicle (V2V) communication is critical to achieving efficient cooperative control. However, in the real-world traffic environment, V2V communication may suffer from time-varying delay and packet loss, leading to degraded control performance and even safety risks. To mitigate the adverse effects of non-ideal communication, this paper proposes a Dynamic Communication Topology based Multi-Agent Reinforcement Learning (DCT-MARL) algorithm for robust cooperative platoon control. Specifically, the state space is augmented with historical control action and delay to enhance robustness against communication delay. To mitigate the impact of packet loss, a multi-key gated communication mechanism is introduced, which dynamically adjusts the communication topology based on the correlation between vehicles and their current communication status. Simulation results demonstrate that the proposed DCT-MARL significantly outperforms state-of-the-art methods in terms of string stability and driving comfort, validating its superior robustness and effectiveness.
@@ I Introduction
In recent years, with the rapid deployment of Vehicle-to-Vehicle (V2V) communication facilities and autonomous driving technologies, connected vehicle platooning has become a promising approach for enhancing road safety and traffic efficiency [1]. Through V2V communication, connected vehicles can form self-organized platoons and maintain a short inter-vehicle distance and coordinated driving behavior, thereby improving road throughput while reducing energy consumption and emissions [2].
@@
In such platooning systems, the effectiveness of cooperative control critically depends on the real-time sharing of vehicle driving state information (e.g., position, velocity, acceleration, etc.). Consequently, reliable V2V communication is an essential prerequisite for ensuring control performance [3]. However, in the real-world traffic environment, channel fading, signal obstruction, and spectral interference may induce substantial communication quality variations, resulting in time-varying delay and packet loss.
@@
These imperfections not only hinder timely and accurate information exchange, as delay leads to outdated data and packet loss results in missing information, but also directly affect the string stability and driving comfort of the platooning system [4]. Hence, developing robust platoon control strategies under non-ideal communication conditions has become a pressing research challenge.
@@
To address these issues, a range of delay compensation based robust control strategies have been proposed. Some studies model deterministic or random delays to improve the robustness of control algorithms [5, 6]. More recently, researchers have also explored solutions within the Reinforcement Learning (RL) framework by expanding the agent’s state space to include historical control actions [7, 8], which improves the system’s robustness under delayed communication. Nevertheless, these methods largely rely on simplified delay assumptions and lack explicit sensitivity to the time-varying nature of delay, limiting their scalability to highly dynamic communication environments.
@@
For packet loss, mitigation strategies typically rely on holding previous inputs, interpolating historical data, or fallback modes [9, 10, 11, 12, 13, 14]. Although these methods can maintain basic control performance under mild packet loss, they often fail to guarantee stability and safety in high-loss or highly dynamic environments. Some recent studies have introduced dynamic communication topology mechanisms that adjust weights based on communication quality, aiming to mitigate performance degradation [15, 16]. While this mitigates the performance degradation from communication packet loss, these strategies focus solely on communication quality metrics and overlook the need to consider the relevance of communication vehicles in the current dynamic traffic environment. In reality, the relevance of each neighboring vehicle is dynamically determined. For example, during acceleration, a driver’s attention naturally shifts to the vehicle ahead, as its motion governs the feasibility of speeding up, whereas during deceleration, the vehicle behind becomes more relevant, since its timely response is critical for avoiding rear-end collisions [17].
@@
Moreover, most existing studies typically address either communication delay or packet loss in isolation, lacking a unified control framework that considers their coupled impact. Although some traditional optimization-based works have proposed integrated modeling approaches [18, 19, 20], they typically rely on accurate system models and analytical formulations, which limits their adaptability and generalization under real-world uncertainties and unmodeled dynamics. Therefore, it’s necessary to investigate novel intelligent algorithms to solve the adaptability problem in dynamic and complex traffic scenarios.
@@
In contrast, Multi-Agent Reinforcement Learning (MARL), has recently emerged as a promising solution for cooperative platoon control due to its model-free nature, strong generalization capabilities, and support for distributed decision making [21, 22, 23]. Under assumptions of ideal communication or deterministic delay, previous studies have validated MARL’s efficacy in platoon control [24, 25]. Some methods further enhance MARL by introducing communication-aware designs, that though consider inter-agent information correlations to build dynamic communication topologies [26]. However, these approaches tend to focus on correlation among agents, while overlooking the dynamic of communication quality.
@@
To bridge this gap, this paper proposes a novel Dynamic Communication Topology based Multi-Agent Reinforcement Learning (DCT-MARL) algorithm to enhance the control performance of connected vehicle platoon under non-ideal communication condition. By dynamically adjusting the communication topology and augmenting the state space, the proposed method improves the robustness and adaptability of vehicle platoon in realistic communication environment. The main contributions are summarized as follows:
@@
    A multi-key gated communication network is proposed, enabling each agent to infer the relevance of communication agents based on real-time observations and historical contexts. This mechanism dynamically adjusts the communication topology in response to packet loss.
@@
    A state enhancement scheme is designed to mitigate the impact of communication delay on platoon control. Specifically, the agent’s state is extended to include both the history control action and delay information. This increase improves the policy sensitivity to time-varying communication delay and improves the robustness and adaptability of the system to delayed communication.
@@
    A co-simulation environment integrating vehicular control and wireless communication is developed to enable comprehensive evaluation under realistic condition. Extensive experiments demonstrate that the proposed DCT-MARL algorithm outperforms state-of-the-art methods in terms of string stability and driving comfort under non-ideal communication condition.
@@ II-A Platoon Control under Unideal Communication Conditions
Reliable V2V communication is essential for cooperative platoon control, however, packet loss and delay in practical V2V communication may induce performance degradation in the control system [3, 4].
@@
To mitigate the negative impact of communication delay, a range of delay compensation based robust control strategies have been proposed [5, 27, 7, 8, 6]. To address packet loss, traditional approaches employ techniques include Zero-Order Hold (ZOH) to maintain previous control inputs [9], historical data interpolation [10, 11], or fallback modes such as Adaptive Cruise Control (ACC) [12, 13, 14]. To enhance adaptability, some recent studies have introduced dynamic communication topology mechanisms that adjust weights based on communication quality, thereby mitigating the effects of communication packet loss [15, 16]. For example, Liu et al. [15] developed an traditional optimization-based method to adjust weights according to communication condition, whereas Wang et al. [16] proposed a self-adaptive topology control method based on a Dual-Decision Deterministic Policy Gradient ($D^{3}PG$) algorithm for platoon control, which significantly enhances system resilience under fluctuating communication condition. However, these strategies primarily focus on communication quality metrics, overlooking the fact that in practical applications, the selection of communication targets should be context-aware and responsive to dynamic traffic conditions.
@@
Furthermore, most existing studies address either delay or packet loss in isolation, making them insufficient for scenarios where both impairments coexist. In real-world environments, however, delay and packet loss frequently occur simultaneously, creating compounded challenges for platoon control systems. To address this, some recent studies have proposed unified models that simultaneously account for the effects of delay and packet loss [5, 9, 18, 19, 20]. For instance, Vegamoor et al. [18] modeled delay and packet loss simultaneously, and introduced a safety-aware constraint on distance errors between any two vehicles and derived a lower bound on safe time gaps to ensure string stability under unideal communication scenarios. Zhao et al. [5] examined control strategies under constant time headway distance, modeling non-ideal communication conditions including limited communication range, random packet loss, and time-varying delay. Their approach employed a distance-based dynamic topology and a robust controller that utilized multi-leader an follower information. Following this, Wang et al. [20] further advanced the approach by proposing a platoon following model based on dynamic weight optimization, which adaptively adjusted the weighting of incoming information. Their approach integrated realistic vehicle dynamics and demonstrated strong robustness under non-ideal communication conditions.
@@
While these control approaches provide valuable theoretical foundations, they often rely heavily on accurate system modeling and incur significant computational overhead. Consequently, they are difficult to scale in dynamic and uncertain traffic environment characterized by rapidly changing velocities and nonlinear interactions.
@@ II-B MARL-based Platoon Control
In recent years, MARL has gained significant attention in intelligent transportation systems due to its strengths in distributed decision-making, policy learning, and environmental adaptability[24, 28, 29, 25]. MARL has become a promising paradigm for cooperative control tasks, such as vehicle platooning. For example, Chen et al.[28] evaluated MARL frameworks using both local and global reward structures in platoons of three and five vehicles, and found that decentralized learning with local rewards achieved superior performance in terms of stability and responsiveness compared to centralized approaches.
However, the vast majority of existing MARL-based platoon control studies assume ideal V2V communication, free of delay and packet loss, and thus lack a comprehensive evaluation into the robustness of learned policies under realistic communication impairments. To address this gap, Wang et al. [25] proposed a delay-aware MARL framework based on a partially observable Markov game (POMG), capturing delay-induced uncertainties in platoon systems. Their simulations demonstrated improved platoon stability under time-varying delays. However, existing works still fall short of simultaneously addressing communication delays and packet losses in a unified MARL framework.
On dynamic communication topologies within MARL, researchers have explored inter-agent information correlations to construct adaptive communication topologies. Seminal algorithms including ITGCNet[30], Gated-ACML[31], I2C[26], and NeurComm[32] employ gating or attention mechanisms to determine when and with whom to communicate, thereby improving learning efficiency. In particular, Chu et al.[32] proposed a learnable protocol that enables agents to formulate distributed policies based on both local and neighbor information, partially mitigating the adverse effects of communication failure. Ding et al. [26] propose I2C, a dynamic communication framework where each agent learns to infer its communication targets from local observations via a prior network trained through causal inference. However, these methods primarily emphasize interagent information relevance while neglecting the dynamic variability of the communication environment, limiting their applicability under real-world communication environment.
To overcome these limitations, this paper proposes a DCT-MARL framework that jointly addresses communication delay and packet loss. The framework aims to substantially enhance the robustness and adaptability of connected vehicle platoon in non-ideal communication environment.
@@ III-A Vehicle Dynamics Model
This paper considers a platoon consisting of $N+1$ vehicles, indexed as 0,1,,n0,1,…,n, where the leading vehicle is indexed as 0, and the following vehicles are indexed from $1$ to $n$. All vehicles communicate through V2V wireless communication to exchange information and coordinate movements. In this study, this paper primarily focus on a linear model of a connected automated vehicle (CAV) system, which has been extensively used in vehicle control. The vehicle dynamics are given by [33].
@@
In this study, this paper adopt the constant time headway policy (CTHP), meaning that the platoon maintains a constant time gap between adjacent vehicles. Each following vehicle $i$ aims to maintain a desired headway distance from its preceding vehicle i1i-1, which is defined as the bumper-to-bumper distance. The ideal inter-vehicle distance between vehicle $i$ and its predecessor i1i-1 is denoted as $d_{r,i,t}$.
@@
Since all vehicles in the platoon are assumed to be identical, $d_{0}$ and $l$ are considered uniform across all vehicles and are therefore omitted in the subsequent derivations.
For any adjacent vehicle pair $(i,j)$ in the communication graph, the tracking error consists of both distance error and velocity error, defined as follows:
@@
$\mathrm{err}_{v_{i,t}}$ $vi,tvj,t,j>i\displaystyle$=
Here, $j<i$ means vehicle $j$ is ahead of vehicle $i$. The distance error $\text{err}_{p,i,t}$ captures the deviation between the actual distance of vehicles $i$ and $j$ and the cumulative desired distance, which is obtained by summing the product of the desired time headway $h$ and the instantaneous velocities of all intermediate vehicles between $i$ and $j$. The velocity error $\text{err}_{v,i,t}$ reflects the relative speed difference between the two vehicles.
Each vehicle $i$ is capable of communicating with other vehicles within its communication range. However, due to potential communication disruptions or packet loss, the communication agent may become unavailable at certain time steps. Additionally, due to network congestion, signal attenuation, or interference, communication may experience a delay $\xi_{i,j}$, meaning that vehicle $i$ at time $t$ receives information from vehicle $j$ that corresponds to time $t-\xi_{i,j}$ instead of real-time data.
@@ III-B Control Objectives of Vehicle Platoon
where $\mathrm{acc}_{i,t}$ is the actual acceleration of vehicle $i$. The bounds $\mathrm{acc}{\min}$ and $\mathrm{acc}{\max}$ represent the allowable acceleration range, while $u_{\min}$ and $u_{\max}$ denote the bounds on control input. These constraint ensure safe and comfortable driving and prevent excessive control efforts.